Given the information in the table,

Current dividend $6.00

Growth Rate in Dividends 2.50%

Required Return on Equity Rs 8.00%

According to the Gordon Growth Model, what is the $ amount of the Capital Gains or Loss between periods 2 and 3 ?

$3.01

$2.94

$2.98

$2.84
Answer #1

Current price = D1 / required rate - growth rate

Current price = [6(1 + 2.5%)] / 0.08 - 0.025

Current price = 6.15 / 0.055

Current price = 111.81818

Year 2 price = Present value (1 + g)n

Year 2 price = 111.81818 (1 + 0.025)2

Year 2 price = 111.81818 * 1.050625

Year 2 price = 117.478975

Year 3 price = 111.81818 (1 + 0.025)3

Year 3 price = 111.81818 * 1.076891

Year 3 price = 120.41595

Capital gain = 120.41595 - 117.478975

Capital gain = $2.94